# Breaking the Performance Ceiling in Complex Reinforcement Learning requires Inference Strategies

### v1

@@ Abstract
Reinforcement learning (RL) systems have countless applications, from energy-grid management to protein design. However, such real-world scenarios are often extremely difficult, combinatorial in nature, and require complex coordination between multiple agents. This level of complexity can cause even state-of-the-art RL systems, trained until convergence, to hit a performance ceiling which they are unable to break out of with zero-shot inference. Meanwhile, many digital or simulation-based applications allow for an inference phase that utilises a specific time and compute budget to explore multiple attempts before outputting a final solution. In this work, we show that such an inference phase employed at execution time, and the choice of a corresponding inference strategy, are key to breaking the performance ceiling observed in complex multi-agent RL problems. Our main result is striking: we can obtain up to a 126% and, on average, a 45% improvement over the previous state-of-the-art across 17 tasks, using only a couple seconds of extra wall-clock time during execution. We also demonstrate promising compute scaling properties, supported by over 60k experiments, making it the largest study on inference strategies for complex RL to date. Our experimental data and code are available at https://sites.google.com/view/inf-marl [^1] A GitHub repository will be released shortly..
@@ 1 Introduction
Learning to solve sequential decision-making tasks is a central challenge in artificial intelligence (AI), with far-reaching applications ranging from energy-grid optimisation (Ahmad et al., 2021) and autonomous logistics (Laterre et al., 2018) to molecular discovery (Olivecrona et al., 2017) and drug design (Popova et al., 2018). Complex sequential real-world problems that cannot be solved by traditional optimisation techniques are inherently complex and require navigating high-dimensional solution spaces. Reinforcement Learning (RL) presents a promising avenue to improve our capacity to find efficient solutions to these problems, but despite stunning progress over the past decade, such as human-level performance in Atari games (Mnih et al., 2015), defeating the world champion in the game of Go (Silver et al., 2016) or aligning AI systems with human preferences (Stiennon et al., 2020), current approaches are facing challenges that prevent their common deployment in most real-world systems (Dulac-Arnold et al., 2020).
@@
A major source of this difficulty lies in the combinatorial nature of many decision-making tasks. As the problem size increases, the space of possible solutions grows exponentially (Karp, 1975). In multi-agent systems, the challenge compounds: agents must coordinate in environments where only partial information is available, the joint action space is combinatorial, and optimal behaviour depends on precise interaction with other agents (Bernstein et al., 2000; Canese et al., 2021). These properties make it fundamentally difficult to rely on the zero-shot performance of a trained policy, even if that policy was optimised to convergence on a representative training distribution. This causes the gap between zero-shot performance and optimality to grow substantially with increasing complexity (see Fig. 1).
@@
However, numerous practical applications are not restricted to producing a single zero-shot solution. Instead, inference is often permitted to take place over a few seconds, minutes or hours, with a given computational resource. Furthermore, models and simulators are often accessible and very efficient (e.g., energy grid management, train scheduling, package delivery, routing, printed circuit board design) and provide either an exact score or a very accurate approximation. In other applications where the gap to reality may be larger (e.g., protein design, robotics), improving the solution under the simulated score can still arguably provide significant improvement towards the real objective (Hayes et al., 2025; Dona et al., 2024; Hundt et al., 2019; Rao et al., 2020).
@@
This opens up an opportunity: rather than relying on a single attempt of the trained policy, the time budget and compute capacity can be leveraged to actively search for better solutions using multiple attempts, following an inference-time strategy. For instance, progressively building a tree of possible solutions, or adapting the policy using outcomes of past attempts. Even straightforward strategies can provide significant performance improvement with low time cost. For instance, generating a large batch of diversified solutions in parallel, using stochastic sampling, rather than a single greedy solution: given a modern GPU, this enables to produce hundreds of solutions, for the same wall-clock time, enabling massive exploration at no time cost. These strategies are rarely emphasized in existing benchmarks (Papoudakis et al., 2021; Mahjoub et al., 2025), and many practitioners invest months of research trying to improve the zero-shot performance of their models on scenarios where only marginal improvement may still be achieved. Whereas they could unlock performance gains from inference-time search, at negligible wall-clock time cost with moderate compute capacity (Fig. 1).
@@
Research in RL for Combinatorial Optimisation (CO) has produced efficient inference strategies (Bello et al., 2016; Hottung et al., 2022; Choo et al., 2022; Chalumeau et al., 2023b), often referred to as active search, or online adaptation methods. However, their empirical study is still limited to a few problems, a narrow range of budget settings (Chalumeau et al., 2024b), and barely no insight on their scaling properties. In the multi-agent case, there is no study on inference strategies for collaborative teams: most work on team adaptation focus on Ad Hoc Teamwork (Mirsky et al., 2022; Wang et al., 2024a; Ruhdorfer et al., 2025), which is adjacent to our objective. Interestingly, most recent studies about the impact of inference strategies come from the Large Language Model (LLM) literature (Snell et al., 2025; Muennighoff et al., 2025; Wu et al., 2025), where the adequate combination of efficient models with inference strategies is currently state-of-the-art (SOTA).
@@
In this work, we formalise and investigate the role of inference strategies in complex decision-making tasks. To capture the full complexity of tasks described above, we formulate our problem setting as a decentralised partially observable Markov decision process (Dec-POMDP) (Kaelbling et al., 1998). This is instead of the typical single-agent MDP used in many RL studies. We make this choice for several reasons: (1) it more accurately maps onto many complex real-world problems of interest, (2) Dec-POMDPs subsume MDPs by being strictly more complex (Bernstein et al., 2000), and (3) because of this, we expect our findings to translate to all simpler problem formulations. Within this setting, we provide a unifying view of popular inference-strategy paradigms, including policy sampling, tree search (Choo et al., 2022), online fine-tuning (Bello et al., 2016; Hottung et al., 2022), and diversity-based search (Chalumeau et al., 2023b). Strikingly, we show that across a wide range of specifically selected difficult RL problems, inference strategies boost performance on average by $45$%, over zero-shot SOTA. Furthermore, in the best of cases, this boost can be as large as $126$%. All of this, using only a couple of seconds of additional execution time.
@@
Our results call for a shift in how RL systems are evaluated and deployed: inference strategies are not a minor post-processing step, but a key performance driver in realistic conditions. This work sets the foundation for a more nuanced view of inference in sequential decision-making and provides the tools to build systems that can scale with compute. All our code and experimental data can be accessed at: https://sites.google.com/view/inf-marl.
@@ 2 Related work
Beyond naive stochastic sampling, several paradigms have been explored to generate the best possible solution using a trained policy checkpoint during inference. Online fine-tuning Bello et al. (2016) retrains all policy parameters with RL using past attempts. Hottung et al. (2022) re-trains only a subset of the policy’s parameters to reduce memory and compute overheads, enabling more attempts for a given inference budget. It also adds a imitation learning term to the RL term, to force exploration close to the best solution found so far. Choo et al. (2022) uses tree search, with simulation guided node estimates under budget constraints, which outperforms Beam Search and Monte Carlo Tree Search (Coulom, 2006). Diversity-based methods: inspired by previous diversity-seeking approaches, like unsupervised skill discovery (Eysenbach et al., 2019; Sharma et al., 2019; Kumar et al., 2020) and quality-diversity (Chalumeau et al., 2023a; Cully and Demiris, 2017), Grinsztajn et al. (2023) introduces an RL objective that trains a population of diverse and specialized policies, efficient for few-shot performance. Chalumeau et al. (2023b) uses this objective and encodes the diversity in a continuous latent space that can be searched at inference-time, introducing the SOTA method compass; meanwhile Hottung et al. (2024) uses a similar approach but with a discrete encoding space.
These works have introduced most of the inference strategies we consider in this paper, but they fall short on three important aspects that we aim to improve: (i) they evaluate inference strategies on benchmarks where over $95$% zero-shot optimality is already achieved, leaving little room for meaningful gains, (ii) these benchmarks rely on domain-specific tricks such as starting points or instance augmentations; and (iii) methods are compared under a unique budget setting, overlooking the fact that relative performance depends on the available compute and time budget. In addition, their ability to scale with compute remains unexplored, despite being a critical property.
@@
There is only limited work on inference strategies for MARL. Most work about search and adaptation within MARL focus on the challenge of Ad Hoc Teamwork (Yourdshahi et al., 2018; Hu et al., 2020; Mirsky et al., 2022; Wang et al., 2024a; Ruhdorfer et al., 2025; Hammond et al., 2025), often in the form of zero-shot coordination, where agents must generalize to new partners at execution time. While these lines of work share some methodological similarities, for instance using diversity-seeking training (Long et al., 2024; Lupu et al., 2021) or adapting through tree search (Yourdshahi et al., 2018), they pursue fundamentally different goals and remain orthogonal. In our work, our focus is primarily on solving difficult and complex industrial optimisation tasks.
@@
Recent advances in LLMs are closely intertwined with the use of inference strategies (Snell et al., 2025; Wei et al., 2022; Wang et al., 2024b), and a growing effort has gone into studying their scaling properties (Muennighoff et al., 2025; Wu et al., 2025). However, the typical inference-time setting is usually different from ours. LLMs have very costly forward passes, and cannot access the exact score of their answers, but can approximate them using a reward model (Ouyang et al., 2022). Most popular strategies for LLMs are designed for few shots, namely sampling and ensembling (e.g., majority voting).
Overall, numerous inference strategies have been proposed in the literature, yet their efficiency under various evaluation settings remains unexamined. Multi-agent RL, despite its inherent complexity, rarely considers inference-time search beyond Ad Hoc Teamwork. Moreover, the broader field of decision-making has not systematically studied how inference strategies scale with compute. Our work aims at filling these gaps by extending the evaluation of inference strategies in RL, demonstrating major performance gains over a wide range of budget settings with impressive scaling properties.
@@ 3.1 Preliminaries
We focus on RL approaches, and use a neural network (policy) that can construct a solution by taking a sequence of actions. This policy is optimised during a training phase and then used during an inference phase, along with an inference strategy, to construct the best possible solution to a new problem instance under a given time and compute budget. These two phases, illustrated in Fig. 2, have different assumptions, objectives and constraints, detailed in the following paragraphs.
@@
We assume that each problem instance can be formulated as a Dec-POMDP (Kaelbling et al., 1998), defined by the tuple $\mathcal{M}=\bigl{(}N,\,\mathcal{S},\,\boldsymbol{\mathcal{O}},\,\boldsymbol{\Omega},\,\boldsymbol{\mathcal{A}},\,R,\,P,\,\gamma,\,H\bigr{)}$. Here $N$ is the number of agents, $\mathcal{S}$ is the environment state space, $\boldsymbol{\mathcal{O}}=\prod_{i=1}^{N}\mathcal{O}^{i}$ is the joint agent observation space, $\boldsymbol{\Omega}:\mathcal{S}\mapsto\boldsymbol{\mathcal{O}}$ is the observation function, $\boldsymbol{\mathcal{A}}=\prod_{i=1}^{N}\mathcal{A}^{i}$ is the joint action space, $R:\mathcal{S}\times\boldsymbol{\mathcal{A}}\mapsto\mathbb{R}$ is the shared reward function, $P:\mathcal{S}\times\boldsymbol{\mathcal{A}}\mapsto\Delta\mathcal{S}$ is the environment transition function, the scalar $\gamma\in[0,1]$ is the discount factor and $H$ is the finite episode horizon. At each timestep the environment occupies a state $s_{t}$ which is mapped to a joint partial observation $\boldsymbol{o}_{t}$ via $\boldsymbol{\Omega}$. The joint action is then sampled following the joint policy, $\boldsymbol{a}_{t}\sim\boldsymbol{\pi}(\,\cdot\mid\boldsymbol{o_{t}})$ and is executed in the environment; after which the environment transitions to the next state $s_{t+1}$, following $P(\,\cdot\mid s_{t},\boldsymbol{a}_{t})$, and the team receives shared the reward $r_{t}=R(s_{t},\boldsymbol{{a}_{t}})$.
@@
We assume a distribution of problem instances $\mathcal{D}$, that can be sampled from. The joint policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}$, parameterised by $\boldsymbol{\theta}$, is used to construct a solution sequentially by taking joint actions conditioned on the joint observation at each timestep. We use RL to train this policy to maximise the expected return obtained when building solutions to instances drawn from the distribution $\mathcal{D}$ over a horizon $H$: $J(\boldsymbol{\pi}_{\boldsymbol{\theta}})=\mathbb{E}_{\mathcal{D}}\left[\sum_{t=0}^{H}\gamma^{t}R(s_{t},\boldsymbol{a_{t}})\right]$. This training objective corresponds to a single attempt (zero-shot). Ideally, this objective should anticipate the multiple attempts allowed at inference, but this is hard to scale. Recent works incorporate such few-shot objectives (Grinsztajn et al., 2023; Chalumeau et al., 2023b, 2024b), but none can yet scale beyond $200$ attempts.
@@
where $\text{C}(\mathcal{I})$ and $\text{T}(\mathcal{I})$ represent the compute and time cost of the inference strategy. This formulation highlights that, unlike traditional RL, where zero-shot performance is the primary measure, we focus on strategies which enable further improvement under given constraints.
3.2 One budget, many uses: inference-time search and adaptation
In this section, we detail four types of inference strategies and how we adapt them to work in the multi-agent setting. We implement and release all of these methods in JAX (Bradbury et al., 2018).

### v3

# Breaking the Performance Ceiling in Reinforcement Learning requires Inference Strategies

@@ Abstract
Reinforcement learning (RL) systems have countless applications, from energy-grid management to protein design. However, such real-world scenarios are often extremely difficult, combinatorial in nature, and require complex coordination between multiple agents. This level of complexity can cause even state-of-the-art RL systems, trained until convergence, to hit a performance ceiling which they are unable to break out of with zero-shot inference. Meanwhile, many digital or simulation-based applications allow for an inference phase that utilises a specific time and compute budget to explore multiple attempts before outputting a final solution. In this work, we show that such an inference phase employed at execution time, and the choice of a corresponding inference strategy, are key to breaking the performance ceiling observed in complex multi-agent RL problems. Our main result is striking: we can obtain up to a 126% and, on average, a 45% improvement over the previous state-of-the-art across 17 tasks, using only a couple seconds of extra wall-clock time during execution. We also demonstrate promising compute scaling properties, supported by over 60k experiments, making it the largest study on inference strategies for complex RL to date. Our experimental data and code are available at https://sites.google.com/view/inference-strategies-rl.
@@ 1 Introduction
Learning to solve sequential decision-making tasks is a central challenge in artificial intelligence (AI), with far-reaching applications ranging from energy-grid optimisation (AHMAD2021125834) and autonomous logistics (laterre2018rankedrewardenablingselfplay) to molecular discovery (olivecrona2017molecular) and drug design (popova2018deep). Complex sequential real-world problems that cannot be solved by traditional optimisation techniques are inherently complex and require navigating high-dimensional solution spaces. Reinforcement Learning (RL) presents a promising avenue to improve our capacity to find efficient solutions to these problems, but despite stunning progress over the past decade, such as human-level performance in Atari games (mnih2015dqn), defeating the world champion in the game of Go (silver2016mastering) or aligning AI systems with human preferences (stiennon2020summarizehf), current approaches are facing challenges that prevent their common deployment in most real-world systems (dulac2020empirical).
@@
A major source of this difficulty lies in the combinatorial nature of many decision-making tasks. As the problem size increases, the space of possible solutions grows exponentially (karp1975computational). In multi-agent systems, the challenge compounds: agents must coordinate in environments where only partial information is available, the joint action space is combinatorial, and optimal behaviour depends on precise interaction with other agents (bernstein2000complexity; canese2021marl). These properties make it fundamentally difficult to rely on the zero-shot performance of a trained policy, even if that policy was optimised to convergence on a representative training distribution. This causes the gap between zero-shot performance and optimality to grow substantially with increasing complexity (see Fig.˜1).
@@
However, numerous practical applications are not restricted to producing a single zero-shot solution. Instead, inference is often permitted to take place over a few seconds, minutes or hours, with a given computational resource. Furthermore, models and simulators are often accessible and very efficient (e.g., energy grid management, train scheduling, package delivery, routing, printed circuit board design) and provide either an exact score or a very accurate approximation. In other applications where the gap to reality may be larger (e.g., protein design, robotics), improving the solution under the simulated score can still arguably provide significant improvement towards the real objective (esm3; dona2024qualitydiversity; hundt2019goodrobot; rao2020rlcyclegan).
@@
This opens up an opportunity: rather than relying on a single attempt of the trained policy, the time budget and compute capacity can be leveraged to actively search for better solutions using multiple attempts, following an inference-time strategy. For instance, progressively building a tree of possible solutions, or adapting the policy using outcomes of past attempts. Even straightforward strategies can provide significant performance improvement with low time cost. For instance, generating a large batch of diversified solutions in parallel, using stochastic sampling, rather than a single greedy solution: given a modern GPU, this enables to produce hundreds of solutions, for the same wall-clock time, enabling massive exploration at no time cost.
These strategies are rarely emphasized in existing benchmarks (papoudakis2021benchmarking; mahjoub2025sable), and many practitioners invest months of research trying to improve the zero-shot performance of their models on scenarios where only marginal improvement may still be achieved. Whereas they could unlock performance gains from inference-time search, at negligible wall-clock time cost with moderate compute capacity (Fig.˜1).
@@
Research in RL for Combinatorial Optimisation (CO) has produced efficient inference strategies (Bello16; hottung2022efficient; choo2022simulationguided; chalumeau2023combinatorial), often referred to as active search, or online adaptation methods. However, their empirical study is still limited to a few problems, a narrow range of budget settings (chalumeau2024memento), and barely no insight on their scaling properties. In the multi-agent case, there is no study on inference strategies for collaborative teams: most work on team adaptation focus on Ad Hoc Teamwork (mirsky2022adhoc; wang2024nagent; ruhdorfer2025overcooked), which is adjacent to our objective. Interestingly, most recent studies about the impact of inference strategies come from the Large Language Model (LLM) literature (snell2025scaling; muennighoff2025s1simpletesttimescaling; wu2025inferencescalinglawsempirical), where the adequate combination of efficient models with inference strategies is currently state-of-the-art (SOTA).
@@
In this work, we formalise and investigate the role of inference strategies in complex decision-making tasks. To capture the full complexity of tasks described above, we formulate our problem setting as a decentralised partially observable Markov decision process (Dec-POMDP) (kaelbling1998planning). This is instead of the typical single-agent MDP used in many RL studies. We make this choice for several reasons: (1) it more accurately maps onto many complex real-world problems of interest, (2) Dec-POMDPs subsume MDPs by being strictly more complex (bernstein2000complexity), and (3) because of this, we expect our findings to translate to all simpler problem formulations. Within this setting, we provide a unifying view of popular inference-strategy paradigms, including policy sampling, tree search (choo2022simulationguided), online fine-tuning (Bello16; hottung2022efficient), and diversity-based search (chalumeau2023combinatorial). Strikingly, we show that across a wide range of specifically selected difficult RL problems, inference strategies boost performance on average by 4545%, over zero-shot SOTA. Furthermore, in the best of cases, this boost can be as large as 126126%. All of this, using only a couple of seconds of additional execution time.
@@
Our results call for a shift in how RL systems are evaluated and deployed: inference strategies are not a minor post-processing step, but a key performance driver in realistic conditions. This work sets the foundation for a more nuanced view of inference in sequential decision-making and provides the tools to build systems that can scale with compute. All our code and experimental data can be accessed at: https://sites.google.com/view/inference-strategies-rl.
@@ 2 Related work
Beyond naive stochastic sampling, several paradigms have been explored to generate the best possible solution using a trained policy checkpoint during inference. Online fine-tuning (Bello16) retrains all policy parameters with RL using past attempts. hottung2022efficient re-trains only a subset of the policy’s parameters to reduce memory and compute overheads, enabling more attempts for a given inference budget, and adds an imitation learning term to the RL term, to force exploration close to the best solution found so far. macfarlane2024spo also investigates inference-time policy improvement via sequential updates. choo2022simulationguided uses tree search, with simulation guided node estimates under budget constraints, which outperforms Beam Search and Monte Carlo Tree Search (coulom2006mcts). Diversity-based methods: inspired by previous diversity-seeking approaches, like unsupervised skill discovery (eysenbach2018diversity; sharma2019dynamics; kumar2020one) and quality-diversity (chalumeau2023neuroevolution; cully2017quality). poppy introduces an RL objective that trains a population of diverse and specialized policies, efficient for few-shot performance. chalumeau2023combinatorial uses this objective and encodes the diversity in a continuous latent space that can be searched at inference-time, introducing the SOTA method compass; meanwhile hottung2024polynet uses a similar approach but with a discrete encoding space.
These works have introduced most of the inference strategies we consider in this paper, but they fall short on three important aspects that we aim to improve: (i) they evaluate inference strategies on benchmarks where over 9595% zero-shot optimality is already achieved, leaving little room for meaningful gains, (ii) these benchmarks rely on domain-specific tricks such as starting points or instance augmentations; and (iii) methods are compared under a unique budget setting, overlooking the fact that relative performance depends on the available compute and time budget. In addition, their ability to scale with compute remains unexplored, despite being a critical property.
Policy adaptation in Meta-RL and Offline-to-Online RL
Meta-RL and offline-to-online RL both design mechanisms for policy adaptation, thereby sharing close conceptual links with several popular inference strategies. compass (chalumeau2023combinatorial) and VariBAD (zintgraf2021varibad) both condition a policy on a latent space and search it to adapt. MEMENTO (chalumeau2024memento) and RL22 (duan2016rl2fastreinforcementlearning) similarly rely on memory and a learned rule to search the policy space. DIMES (qiu2022dimes) explicitly performs meta-RL following MAML’s methodology (finn2017maml). While we are not aware of inference strategies directly inspired by the offline-to-online RL literature, nakamoto2023calql and mark2025policyagnostic introduce mechanisms which could be beneficial at inference time.
It is worth highlighting fundamental distinctions between these fields. Inference strategies focus on (i) the solutions found rather than the learned policy, (ii) maximum rather than average performance, and (iii) single instances at test time, whereas most meta-RL and offline-to-online RL methods adapt from a distribution to another distribution to improve generalisation.
@@
There is only limited work on inference strategies for MARL. Most work about search and adaptation within MARL focus on the challenge of Ad Hoc Teamwork (yourdshahi2018largescaleadhoc; hu2020otherplay; mirsky2022adhoc; wang2024nagent; ruhdorfer2025overcooked; hammond2025symmetrybreaking), often in the form of zero-shot coordination, where agents must generalize to new partners at execution time. While these lines of work share some methodological similarities, for instance using diversity-seeking training (long2024roleplaylearningadaptive; lupu2021trajectory) or adapting through tree search (yourdshahi2018largescaleadhoc), they pursue fundamentally different goals and remain orthogonal. In our work, our focus is primarily on solving difficult and complex industrial optimisation tasks.
@@
Recent advances in LLMs are closely intertwined with the use of inference strategies (snell2025scaling; wei2022chainofthoughts; wang2024hypothesis), and a growing effort has gone into studying their scaling properties (muennighoff2025s1simpletesttimescaling; wu2025inferencescalinglawsempirical). However, the typical inference-time setting is usually different from ours. LLMs have very costly forward passes, and cannot access the exact score of their answers, but can approximate them using a reward model (ouyang2022rlhf). Most popular strategies for LLMs are designed for few shots, namely sampling and ensembling (e.g., majority voting).
Overall, numerous efficient inference strategies have been proposed in the literature, yet their efficiency under various settings remains unexamined. Multi-agent RL, despite its inherent complexity, rarely considers inference-time search beyond Ad Hoc Teamwork. Moreover, the broader field of decision-making has not systematically studied how inference strategies scale with compute. Our work aims at filling these gaps by extending the evaluation of inference strategies in RL, demonstrating major performance gains over a wide range of budget settings with impressive scaling properties.
@@ 3.1 Preliminaries
We focus on RL approaches, and use a neural network (policy) that can construct a solution by taking a sequence of actions. This policy is optimised during a training phase and then used during an inference phase, along with an inference strategy, to construct the best possible solution to a new problem instance under a given time and compute budget. These two phases, illustrated in Fig.˜2, have different assumptions, objectives and constraints, detailed in the following paragraphs.
@@
We assume that each problem instance can be formulated as a Dec-POMDP (kaelbling1998planning), defined by the tuple $\mathcal{M}=\bigl(N,\,\mathcal{S},\,\boldsymbol{\mathcal{O}},\,\boldsymbol{\Omega},\,\boldsymbol{\mathcal{A}},\,R,\,P,\,\gamma,\,H\bigr)$. Here $N$ is the number of agents, $\mathcal{S}$ the environment state space, $\boldsymbol{\mathcal{O}}=\prod_{i=1}^{N}\mathcal{O}^{i}$ the joint agent observation space, $\boldsymbol{\Omega}:\mathcal{S}\mapsto\boldsymbol{\mathcal{O}}$ the observation function, $\boldsymbol{\mathcal{A}}=\prod_{i=1}^{N}\mathcal{A}^{i}$ the joint action space, $R:\mathcal{S}\times\boldsymbol{\mathcal{A}}\mapsto\mathbb{R}$ the shared reward function, $P:\mathcal{S}\times\boldsymbol{\mathcal{A}}\mapsto\Delta\mathcal{S}$ the environment transition function, the scalar $\gamma\in[0,1]$ is the discount factor and $H$ the finite episode horizon. At each timestep, the environment occupies a state $s_{t}$ which is mapped to a joint partial observation $\boldsymbol{o}_{t}$ via $\boldsymbol{\Omega}$. The joint action is then sampled following the joint policy, $\boldsymbol{a}_{t}\sim\boldsymbol{\pi}(\,\cdot\mid\boldsymbol{o_{t}})$ and is executed; after which the environment transitions to the next state $s_{t+1}$, following $P(\,\cdot\mid s_{t},\boldsymbol{a}_{t})$, and the team receives a reward $r_{t}=R(s_{t},\boldsymbol{{a}_{t}})$.
@@
We assume a distribution of problem instances $\mathcal{D}$, that can be sampled from during training. The joint policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}$, parameterised by $\boldsymbol{\theta}$, is used to construct a solution sequentially by taking joint actions conditioned on the joint observation at each timestep. We use RL to train this policy to maximise the expected return obtained when building solutions to instances drawn from the distribution $\mathcal{D}$ over a horizon $H$: $J(\boldsymbol{\pi}_{\boldsymbol{\theta}})=\mathbb{E}_{\mathcal{D}}\left[\sum_{t=0}^{H}\gamma^{t}R(s_{t},\boldsymbol{a_{t}})\right]$.
This training objective corresponds to a single attempt (zero-shot). Ideally, this objective should anticipate the multiple attempts allowed at inference, but this is hard to scale. Recent works incorporate such few-shot objectives (poppy; chalumeau2023combinatorial; chalumeau2024memento), but none can yet scale beyond $200$ attempts.
@@
where $\text{C}(\mathcal{I})$ and $\text{T}(\mathcal{I})$ represent the compute and time cost of the inference strategy. This formulation highlights that, unlike traditional RL, where zero-shot performance is the primary measure, we focus on strategies which enable further improvement under given constraints. We provide three real-world examples in Appendix˜H that correspond to this problem setting, illustrating practical scenarios where inference-time search is both natural and feasible.
Inference strategies differ in how they explore the solution space and how they incorporate the outcomes of previous attempts to influence future sampling. Their effectiveness depends on several contextual factors, including the parameter count of the pre-trained policy, the problem’s underlying structure, the episode horizon, the nature of the reward function, but most critically, on the available time and compute budget.
3.2 One budget, many possibilities: inference-time search and adaptation
In this section, we detail four types of inference strategies and how we adapt them to work in the multi-agent setting. We implement and release all of these methods in JAX (jax2018github).